A bus starting to from rest moves with a uniform acceleration of 0.1m/s for 2 minutes. Find the speed acquired and distance travelled
1
Expert's answer
2020-05-13T09:48:38-0400
"v=at=(0.1)(120)=12\frac{m}{s}"

"s=0.5at^2=0.5(0.1)(120)^2=720\ m"